# Adversarial Pruning: A Survey and Benchmark of Pruning Methods for Adversarial Robustness

### v1

@@ Abstract
Recent work has proposed neural network pruning techniques to reduce the size of a network while preserving robustness against adversarial examples, i.e., well-crafted inputs inducing a misclassification. These methods, which we refer to as adversarial pruning methods, involve complex and articulated designs, making it difficult to analyze the differences and establish a fair and accurate comparison. In this work, we overcome these issues by surveying current adversarial pruning methods and proposing a novel taxonomy to categorize them based on two main dimensions: the pipeline, defining when to prune; and the specifics, defining how to prune. We then highlight the limitations of current empirical analyses and propose a novel, fair evaluation benchmark to address them. We finally conduct an empirical re-evaluation of current adversarial pruning methods and discuss the results, highlighting the shared traits of top-performing adversarial pruning methods, as well as common issues. We welcome contributions in our publicly-available benchmark at https://github.com/pralab/AdversarialPruningBenchmark.
Keywords Adversarial Machine Learning  $\cdot$ Neural Network Pruning
@@ 1 Introduction
Deep neural networks tend to be trained in over-parameterized regimes, where the number of model parameters exceeds the dimension of the training set [1]. Although the performance reached corroborates the need for such a regime, there is still a consistent search for models that suit a resource-constrained scenario, one where the models’ dimensions cannot be chosen at will but rather minimized. In addition, it has been shown that although most of these parameters are fundamental to finding a good minimum during training, they then become superfluous [2] It has thus become of great interest to study techniques capable of reducing the network size while still being able to generalize well. Compression methods such as pruning [3], quantization [4] and knowledge distillation [5], have been therefore studied and adapted to current networks. In this regard, one popular compression approach is neural network pruning, which originates in the late 80s [6] and flourishes with the rise of deep learning [3, 7, 8, 9]. The goal of pruning is to reduce the number of network parameters, i.e. the network size, with as little performance degradation as possible. However, when deployed into security-critical scenarios, neural networks might also be required to be robust against adversarial attacks. In particular, machine-learning models have been discovered to be susceptible to adversarial examples, i.e., small perturbations added to the input data that can mislead models’ predictions [10, 11]. Following the discovery of this phenomenon, an intensive line of research has focused on designing both adversarial attacks [12, 13], as well as defenses [12, 14]. The robustness of the models against such attacks becomes highly relevant in a wide variety of security-related applications, ranging from self-driving cars to modern cybersecurity tasks [15]. In this regard, when ML models need to be both robust against adversarial attacks and compressed, Adversarial Pruning (AP) methods stand out by precisely fulfilling this dual need: creating a pruned network with a given sparsity while preserving robustness. Over the last years, various designs have been used to generate AP methods: while some work follow the conventional three-staged pipeline approach proposed in [3], by pruning a dense model after pretraining and then finetuning it [16, 17], others prune during [18] or before [19] training; while some work optimize a score assigned to each parameter to decide how to prune [16, 20], others aim to solve a constrained optimization problem [17, 21]. These methods can thus be broadly different in terms of design. However, such diversity is hard to contemplate: in fact, the lack of a general framework to categorize State-of-the-Art (SoA) AP methods makes the literature complex and unable to be described in a clear and systematic way. Consequently, when a new AP method is compared with the SoA, the design differences between the competing methods tend to be not considered, ultimately limiting the understanding of each AP design.
In this work, we survey current AP methods and propose a novel taxonomy to categorize them. We find the pruning pipeline and the pruning specifics to be the two major categories under which the design of an AP method can be described. While the pipeline describes when the method prunes the network (e.g., after or during training), the specifics describe how the parameters are removed (e.g., the criterion based on which a parameter is deemed unimportant and pruned). We list punctual details for these two pillars, ultimately outlining each AP method as a whole. Then, we extend our analysis of previous work by analyzing the adversarial robustness evaluation, highlighting the main limitations of SoA AP methods which we often find to be below par with respect to recent advances [22, 23]. To address this issue, we propose a novel unified re-evaluation benchmark, where we test the AP methods in a comparable way and with SoA adversarial attacks, thus yielding a fair and reliable comparison. Being AP methods tested against different experimental setups, it is difficult to rank them and choose the ones that lead to the best models. Our benchmark enables a direct comparison of the methods and their effects on the model’s robustness. We finally re-evaluate the SoA AP methods and analyze, given the taxonomy, the effect of the different designs. With our taxonomy and benchmark, jointly represented in Figure 1, our goal is to foster AP methods to be categorically described and fairly tested, thus serving as a “blueprint" for newly published APs.
Figure 1: Overall of the contributions. We propose a taxonomy of Adversarial Pruning (AP) methods, based on the pipeline and specifics, through which we are able to classify the AP methods. Then, via the proposed benchmark, we define a clear experimental setup and reliable adversarial setting through which we enable a fair comparison of the AP methods.
@@ 2 Background
This section introduces the basic concepts behind neural network pruning and adversarial robustness. The goal is to provide a clear understanding of pruning and how it can be designed to create sparse networks capable of preserving robustness against adversarial examples.
@@ 2.1 Neural Network Pruning
Like many other modern techniques applied to deep learning, neural network pruning originated in the late 80s [6, 24]. With the rise of deep learning, networks became computationally intensive and memory-requiring, thus often trespassing resource-constrained scenarios where the available memory and computational capability might not be enough to make modern large networks work [25]. Consequently, the research community became increasingly interested in neural network pruning techniques, which found parameters to be fundamental during training and excessive afterwards, evolving to highly efficient methods [3, 7, 9].
2.1.1 General Pruning Formulation
The goal of pruning is to reduce network size, by removing single weights or entire structures (e.g., filters), while not jeopardizing performances. In this formulation, we collectively refer to single weights, entire filters, channels, or kernels as “parameters". The pruning problem starts from a desired sparsity rate $s_{r}$, and removes enough parameters to satisfy it (i.e., sets them as equal to 00). Therefore, pruning amounts to create a mask $\bm{m}^{*}$ formalized as follows:
$\bm{m}^{*}=\underset{\|\bm{m}\|_{0}\leq k}{\arg\min}$ $\mathcal{L}(\bm{\theta}\otimes\bm{m},\bm{x},y)$ (1)
where $\bm{\theta}$ represents the set of parameters learned by minimizing loss $\mathcal{L}$, and $(\bm{x},y)\sim\mathcal{D}$. The rate $k$, limiting the $\ell_{0}$ norm of the parameters, indicates the desired limit on the number of parameters to be pruned. For the rest of this work, we will control such quantity through a sparsity rate, $s_{r}$ $\in[0,1]$, defined as $s_{r}=1-k/|\bm{\theta}|$. The mask $\bm{m}$, can be instead described as:
$\bm{m}=\begin{cases}0&\text{ if }\theta_{i}\in\mathcal{C}\\ 1&\text{ otherwise }\end{cases}\text{ , }\forall\text{ }\theta_{i}\in\bm{\theta}\end{aligned}\,,$ (2)
@@ 2.1.1 General Pruning Formulation
where the set $\mathcal{C}$ indicates whether a parameter $\theta_{i}\in\bm{\theta}$ is deemed to be pruned. The optimization problem in Eq. 1, being the mask $\bm{m}$ multiplied to the weights binary as defined in Eq. 2, represents a non-convex, combinatorial problem, thus not directly solvable through common approaches such as Stochastic Gradient Descent (SGD). For this reason, recent pruning methods have developed different specifics, defining how to prune, as well as different pipelines, defining when to prune. In terms of specifics, multiple approaches have circumvented the problem of Eq. 1 by using optimizable continuous masks [16], relaxing the problem constraints [17], shrinking the parameters via regularization approaches [9], or directly pruning the parameters according to naive heuristics, such as pruning the weights with the lowest magnitude (LWM) [3]. In terms of pipelines, different methods have proposed likewise different instants of the standard training procedure in which to prune. In this regard, a classical procedure is represented by the three-staged approach proposed in [3], where the network is pretrained, the pruning mask is applied (after training), and the pruned model is then finetuned to restore lost performances. Different approaches instead, prune during training by jointly creating a mask and updating the network parameters [26], or prune before training by finding a sparse subnetwork to be trained from scratch [2].
@@ 2.2 Adversarial Robustness
Reliable adversarial evaluation. Evaluating the adversarial robustness with a single attack, however, is globally considered to represent a not highly informative and often unreliable evaluation [22]. In addition, such attacks are typically run with a default hyperparameter setting, which does not necessarily represent the best attack setting [28]. Therefore, the recent reference method to evaluate adversarial robustness is represented by AutoAttack, which is a parameter-free ensemble of four different attacks typically improving over single attacks such as PGD [13]. Similarly, HO-FMN represents an improved version of the FMN attack algorithm by optimizing the attack hyperparameters [28]. While both attacks enhance the reliability of adversarial evaluations, HO-FMN additionally guarantees to plot security curves efficiently. Security curves depict the robustness decrease of a model against an increasing perturbation budget. However, while in maximum-confidence attacks (or ensembles, such as AA) a single attack run is bounded by a fixed perturbation budget, in minimum-norm attacks the found adversarial examples have a varying perturbation budget which is found by the attack algorithm, thus requiring a single run to plot the curve (while multiple runs would be required for maximum-confidence attacks).
Defenses. Conversely, the goal of a defense is to create a robust model against adversarial attacks. The go-to technique, in this case, is represented by Adversarial Training (AT) [12], whose formulation can be summarized through a min-max optimization problem as follows:
$\min_{\bm{\theta}}\underset{(\bm{x},\bm{y})\sim\mathcal{D}}{\mathbb{E}}\left[\max_{\|\bm{\delta}\|_{p}\leq\epsilon}\mathcal{L}(\bm{\theta},\bm{x}+\bm{\delta},\bm{y})\right]$ (3)
where the perturbation $\delta$ is bounded with the $\epsilon$ value on the $\ell_{p}$ norm, and $(\bm{x},y)\sim\mathcal{D}$ refers to the sampled training data. The inner maximization finds worst-case adversarial perturbation and is typically solved using adversarial attacks such as PGD [12]. The outer problem instead, minimizes the robust loss of the model (i.e., the loss computed on adversarial examples) via SGD, similar to regular training.
@@ 3 Adversarial Pruning
Following the rise of pruning techniques [29], the pursuit of robust models led the community to question the effect of pruning on adversarial robustness [30, 31, 32, 33], thus generating a new spark of research on the interplay between pruning and robustness. We thus surveyed more than 50 papers dealing with pruning and robustness and identified a total of 26 proposing novel approaches to create pruned and robust models. We labelled these methods as Adversarial Pruning (AP) methods, whose goal is to prune the model to a given $s_{r}$ while preserving as much robustness as possible. Initially, AP methods extended the standard three-step pipeline with robust objectives [33] and used a naive pruning criterion (e.g., LWM). However, the research community proposed increasingly complex and diverse designs, employing different pipelines and ways to prune the models while circumventing the problem of Eq. 1. We, therefore, developed a taxonomy of AP methods, presented in Table 1, to systematically classify the APs and have a better understanding of the different designs.
Table 1: The taxonomy of AP methods. We categorize every AP based on pipeline and specifics, and we name the APs with their name/acronym used in the paper. Depending on the pipeline, we subdivide the AP methods into 3 major categories: respectively, methods pruning after (A-methods), before (B-methods) and during (D-methods) training. For each of the three categories, we extend the pipeline with the objective used and the other details, where AT stands for Adversarial Training, NT stands for Natural Training and n.s. stands for Not Specified (hence, not specifying the objective). We then uniformly identify for each AP method a well-defined set of specifics describing how the pruning is applied, indicating whether Structured (S), Unstructured (US), Local (L), or Global (G) pruning are enabled. Finally, the criterion identifies the rule based on which a mask is created.
@@
A-METHODS PRUNING AFTER TRAINING (3.2.1)
@@
B-METHODS PRUNING BEFORE TRAINING (3.2.2)
@@
D-METHODS PRUNING DURING TRAINING (3.2.3)
@@ 3.1 Adversarial Pruning Taxonomy
We defined two main pillars through which AP methods can be classified, pruning pipeline and pruning specifics. The two pillars answer two questions of paramount importance: respectively, when to prune and how to prune. The pipeline, described in Sect. 3.2, precisely indicates the time at which the pruning mask is applied with respect to the training stage. Depending on the pipeline, we identified three major categories of AP methods: pruning after training (A-methods), pruning before training (B-methods), and pruning during training (D-methods). Based on this major categorization, each method defines a set of objectives and details employed in the pipeline. The specifics instead, described in Sect. 3.3, define how the pruning is applied and serve as a scheme for the mask design. We identify, as specifics, the pruning structure, locality, and the pruning criteria.
@@ 3.2.1 A-Methods: Pruning after Training
A-methods, follow the conventional three-step pipeline described in [3], yet extended with robust objectives such as Adversarial Training (AT). Hence, models are adversarially pretrained, pruned, and finetuned with a further robust objective. In this pipeline configuration, the mask $\bm{m}$ is therefore applied to the dense model after the training step.
Pretraining and finetuning. The extension of the pipeline in [3] to adversarial robustness simply requires using robust training objectives. Therefore, within A-methods, we define both pretraining and finetuning objectives. Unsurprisingly, given its wide use, we find PGD AT [12] to be the most commonly employed pretraining and finetuning objective. Every AP method following such pipeline uses AT for both pretraining and finetuning, except for the cases of SR-GKP [49], where natural training (NT) is used for both steps, and PwoA [21], where the pruned model in the finetuning step distils knowledge from the teacher pretrained robust model.

### v2

@@ Abstract
Recent work has proposed neural network pruning techniques to reduce the size of a network while preserving robustness against adversarial examples, i.e., well-crafted inputs inducing a misclassification. These methods, which we refer to as adversarial pruning methods, involve complex and articulated designs, making it difficult to analyze the differences and establish a fair and accurate comparison. In this work, we overcome these issues by surveying current adversarial pruning methods and proposing a novel robustness-oriented taxonomy to categorize them based on two main dimensions: the pipeline, defining when to prune; and the specifics, defining how to prune. We then highlight the limitations of current empirical analyses and propose a novel, fair evaluation benchmark to address them. We finally conduct an empirical re-evaluation of current adversarial pruning methods and discuss the results, highlighting the shared traits of top-performing adversarial pruning methods, as well as common issues. We welcome contributions in our publicly-available benchmark at https://github.com/pralab/AdversarialPruningBenchmark.
keywords:
adversarial machine learning , neural network pruning
journal: Pattern Recognition
[…]
@@ 1 Introduction
Deep neural networks tend to be trained in over-parameterized regimes, where the number of model parameters exceeds the dimension of the training set [1]. Although the performance reached corroborates the need for such a regime, there is still a consistent search for models that suit a resource-constrained scenario, one where the model size cannot be chosen at will but rather minimized. In addition, it has been shown that although most of these parameters are fundamental to finding a good minimum during training, they then become superfluous [2]. It has thus become of great interest to study techniques capable of reducing the network size while still being able to generalize well. Compression methods such as pruning [3], quantization [4] and knowledge distillation [5], have been therefore studied and adapted to current networks. In this regard, one popular compression approach is neural network pruning, which originates in the late 80s [6] and flourishes with the rise of deep learning [3, 7, 8, 9]. The goal of pruning is to reduce the number of network parameters, i.e. the network size, with as little performance degradation as possible.
However, when deployed into security-critical scenarios, neural networks might also be required to be robust against adversarial attacks. In particular, machine-learning models have been discovered to be susceptible to adversarial examples, i.e., small perturbations added to the input data that can mislead models’ predictions [10, 11]. Following the discovery of this phenomenon, an intensive line of research has focused on designing both adversarial attacks [12, 13] and defenses [12, 14] (Sect. 2). The robustness of the models against such attacks becomes highly relevant in a wide variety of security-related applications, ranging from self-driving cars to cybersecurity tasks [15].
In this regard, when Machine Learning (ML) models need to be both robust against adversarial attacks and compressed, Adversarial Pruning (AP) methods stand out by precisely fulfilling this dual need, i.e., creating a pruned network with a given sparsity while preserving robustness. Over the last years, various designs have been used to generate AP methods. In particular, a line of work follows the conventional three-staged pipeline approach proposed in [3], by pruning a robust dense model after pretraining and then finetuning it [16, 17]; other approaches amount to prune the model either during [18] or before [19] training; and further work optimizes a robustness-related score assigned to each parameter to decide how to prune [16, 20], or solve a robust constrained optimization problem directly [17, 21]. These methods can thus be broadly different in terms of design. However, such diversity is hard to contemplate; in fact, the lack of a general framework to categorize state-of-the-art AP methods makes the literature complex and unable to be described in a clear and systematic way. Thus, when a new AP method is compared with competing approaches, their design differences tend to be not properly considered, limiting the understanding of each AP design.
In this work, we survey current AP methods and categorize them within a common taxonomy (Sect. 3). We find the pruning pipeline and the pruning specifics to be the two main axes under which the design of an AP method can be described. While the pipeline describes when the method prunes the network (e.g., after or during training), the specifics describe how the parameters are removed (e.g., the criterion that defines whether a parameter has to be pruned). By detailing the characteristics of each of these axes, we can provide a thorough yet compact description of all AP methods. To the best of our knowledge, we are the first to provide such a categorization for AP methods.
Our work, however, is not limited to providing such a novel categorization. As such methods have been tested using different experimental setups and adversarial attacks, yielding different estimates of their accuracy and adversarial robustness, it is difficult to directly compare them and choose the ones that may lead to the best models. To overcome this issue, in this work we design a unified benchmark to re-evaluate the accuracy and adversarial robustness of current AP methods under the same conditions, and using more recent and reliable attack algorithms. We thus leverage such benchmark to re-evaluate the existing AP methods and analyze, based on our categorization, the effect of their different designs (Sect. 4). With our taxonomy and benchmark, we aim to foster AP methods to be categorically described and fairly tested, thus serving as a “blueprint” for newly published methods. We conclude by discussing related work (Sect. 5), our contributions, and promising future research directions (Sect. 6).
Figure 1: An overview of our taxonomy (left) and our benchmark (right).
@@ 2 Background
This section introduces the basic concepts behind neural network pruning and adversarial robustness. The goal is to provide a clear understanding of pruning and how it can be designed to create sparse networks capable of preserving adversarial robustness.
@@ 2.1 Neural Network Pruning
Like many other modern techniques applied to deep learning, neural network pruning originated in the late 80s [6]. With the rise of deep learning, networks became computationally intensive and memory-requiring, thus often trespassing resource-constrained scenarios where the available memory and computational capability might not be enough to make modern large networks work [22]. Consequently, the research community became increasingly interested in neural network pruning techniques, which found parameters to be fundamental during training and excessive afterward, evolving to highly efficient methods [3, 7, 9].
Pruning formulation. The goal of pruning is to reduce network size, by removing single weights or entire structures (e.g., filters), while not jeopardizing the performance. In this formulation, we collectively refer to single weights, entire filters, channels, or kernels as “parameters,” and use the symbol $\bm{\theta}\in\mathbb{R}^{p}$ to denote them, with $p$ being their overall number. Let us denote with $\mathcal{D}=(\bm{x}_{i},y_{i})_{i=1}^{n}$ the training set, consisting of $n$ $d$-dimensional samples $\bm{x}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, [^1] When data is normalized, as for images, typically $\bm{x}\in\mathcal{X}=[0,1]^{d}$. along with their class labels $y_{i}\in\mathcal{Y}=\{1,\ldots,c\}$. Training an ML model amounts to finding a function (within a feasible set, e.g., constrained by a fixed network architecture) that minimizes a loss $\mathcal{L}(\bm{\theta},\bm{x},y)$ on the training samples in $\mathcal{D}$. This typically requires using a convex loss that provides an upper bound on the classification error (i.e., the zero-one loss), while penalizing complex solutions via a regularization term that prevents overfitting. The pruning problem starts from a desired sparsity rate $s_{r}\in[0,1]$, which amounts to retaining only $k=\lfloor p\cdot(1-s_{r})\rfloor$ non-zero parameters. We can then formalize it as the problem of finding a binary mask $\bm{m}\in\{0,1\}^{p}$ that only retains $k$ parameters while still minimizing the given loss function:
$\bm{m}^{*}\in\underset{\|\bm{m}\|_{0}\leq k}{\arg\min}$ $\sum_{i=1}^{n}\mathcal{L}(\bm{\theta}\odot\bm{m},\bm{x}_{i},y_{i})\,,$ (1)
being $\|\bm{m}\|_{0}$ the $\ell_{0}$ norm of the mask, counting its non-zero elements, and $\odot$ the element-wise (Hadamard) product. We will refer to the algorithms used to solve this problem as $\mathcal{C}$, being represented by the criterion used to select which parameters to prune or not.
@@
The optimization problem in Eq. 1 represents a non-convex, combinatorial problem, thus not directly solvable through common approaches such as Stochastic Gradient Descent (SGD). For this reason, recent pruning methods have developed different specifics, defining how to prune, as well as different pipelines, defining when to prune. In terms of specifics, multiple approaches have circumvented the problem of Eq. 1 by using optimizable continuous masks [16], relaxing the problem constraints [17], shrinking the parameters via regularization approaches [9], or directly pruning the parameters according to naive heuristics, such as pruning the weights with the lowest magnitude (LWM) [3]. In terms of pipelines, different methods have proposed likewise different instants of the standard training procedure in which to prune. In this regard, a classical procedure is represented by the three-staged approach proposed in [3], where the network is pretrained, the pruning mask is applied (after training), and the pruned model is then finetuned to restore its performance. Different approaches, instead, prune during training by jointly creating a mask and updating the network parameters [23], or prune before training by finding a sparse subnetwork to be trained from scratch [2].
@@ 2.2 Adversarial Robustness
Reliable adversarial evaluation. Evaluating the adversarial robustness with a single attack, however, is globally considered to represent a not highly informative and often unreliable evaluation [25]. In addition, such attacks are typically run with a default hyperparameter setting, which does not necessarily represent the best attack setting [26]. Therefore, the recent reference method to evaluate adversarial robustness is represented by AutoAttack, which is a parameter-free ensemble of four different attacks typically improving over single attacks such as PGD [13]. Similarly, HO-FMN represents an improved version of the FMN attack algorithm by optimizing the attack hyperparameters [26]. While both attacks enhance the reliability of adversarial evaluations, HO-FMN additionally guarantees to plot robustness curves efficiently. Robustness curves depict the robustness decrease of a model against an increasing perturbation budget. However, while in maximum-confidence attacks (or ensembles, such as AA), a single attack run is bounded by a fixed perturbation budget, in minimum-norm attacks the found adversarial examples have a varying perturbation budget which is found by the attack algorithm, thus requiring a single run to plot the curve (while multiple runs would be required for maximum-confidence attacks).
Defenses. Conversely, the goal of a defense is to create a robust model against adversarial attacks. The go-to technique, in this case, is represented by Adversarial Training (AT) [12], which is formalized as a min-max optimization problem:
$\min_{\bm{\theta}}\sum_{i=1}^{n}\max_{\|\bm{\delta}_{i}\|_{p}\leq\epsilon}\mathcal{L}(\bm{\theta},\bm{x}_{i}+\bm{\delta}_{i},y)\,,$ (2)
where the $\ell_{p}$ norm of the perturbation $\bm{\delta}$ is upper bounded by $\epsilon$. Note also that $\bm{x}+\bm{\delta}\in\mathcal{X}$, i.e., the perturbed sample has to belong to the input space $\mathcal{X}$. The inner maximization finds the worst-case adversarial perturbation, and it is typically solved using adversarial attacks such as PGD [12]. The outer problem, instead, minimizes the robust loss (computed on the adversarial examples) via SGD, as normally done during training.
@@ 3 Adversarial Pruning
Following the rise of pruning techniques [27, 28], the pursuit of robust models led the community to question the effect of pruning on adversarial robustness [29, 30, 31, 32], thus generating a new spark of research on the interplay between pruning and robustness. We thus surveyed more than 50 papers dealing with pruning and robustness and identified a total of 26 proposing novel approaches to create pruned and robust models. We labeled these methods as Adversarial Pruning (AP) methods, whose goal is to prune the model to a given $s_{r}$ while preserving as much robustness as possible.
Initially, AP methods extended standard three-step pipeline with robust objectives [32] and used a naive pruning criterion (e.g., LWM). However, the research community proposed increasingly complex and diverse designs, employing different robustness-oriented approaches to prune the models while circumventing the problem of Eq. 1. We, therefore, developed a taxonomy of AP methods, presented in Table 1, to systematically classify the APs and have a better understanding of the different designs.
Table 1: The taxonomy of AP methods. We categorize every AP based on pipeline and specifics, and we name the APs with their name/acronym used in the paper. Depending on the pipeline, we subdivide the AP methods into 3 major categories: respectively, methods pruning after (A-methods), before (B-methods), and during (D-methods) training. For each of the three categories, we extend the pipeline with the objective used and the other details, where AT stands for Adversarial Training, NT stands for Natural Training, and n.s. stands for Not Specified (hence, not specifying the objective). We then uniformly identify for each AP method a well-defined set of specifics describing how the pruning is applied, indicating whether Structured (S), Unstructured (US), Local (L), or Global (G) pruning is enabled. Finally, the criterion identifies the rule based on which a mask is created.
@@
      A-Methods: Pruning after Training (3.2.1)
@@
      B-Methods: Pruning before Training (3.2.2)
@@
      D-Methods: Pruning during Training (3.2.3)
@@ 3.1 Adversarial Pruning Taxonomy
We defined two main pillars through which AP methods can be classified, pruning pipeline and pruning specifics. The two pillars answer two questions of paramount importance: respectively, when to prune and how to prune. The pipeline, described in Sect. 3.2, precisely indicates the time at which the pruning mask is applied with respect to the training stage. Depending on the pipeline, we identified three major categories of AP methods: pruning after training (A-methods), pruning before training (B-methods), and pruning during training (D-methods). Based on this major categorization, each method defines a set of robust objectives and details employed in the pipeline. The specifics instead, described in Sect. 3.3, define how the pruning is applied and serve as a scheme for the mask design. We identify, as specifics, the pruning structure, locality, and the pruning criteria.
@@ 3.2.1 A-Methods: Pruning after Training
A-methods, follow the conventional three-step pipeline described in [3] extended with robust objectives such as Adversarial Training (AT). Hence, models are pretrained, pruned, and finetuned with robust objectives. In this pipeline configuration, the mask $\bm{m}$ is therefore applied to the dense model after the training step.
Pretraining and finetuning. The extension of the pipeline in [3] to adversarial robustness simply requires using robust training objectives. Therefore, within A-methods, we define both pretraining and finetuning objectives. Unsurprisingly, given its wide use, we find PGD AT [12] to be the most commonly employed pretraining and finetuning objective. Every AP method following such pipeline uses AT for both pretraining and finetuning, except for the cases of SR-GKP [48], where natural training (NT) is used for both steps, and PwoA [21], where the pruned model in the finetuning step distills knowledge from the teacher pretrained robust model.